A tree casts a shadow 6 feet long. At the same time of day a 5.5 foot tall person casts a shadow 3 feet tall. How tall is the tree
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

x/6=5.5/3
3x=6*5.5
x=33/3
x=11ft

Answer by stanbon(75887)

A tree casts a shadow 6 feet long. At the same time of day a 5.5 foot tall person casts a shadow 3 feet tall. How tall is the tree.
---------
object/shadow = object/shadow
--------------------------------
tree/6 = 5.5/3
-------------------
tree = 6(5.5/3) = 2(5.5) = 11 ft
-----------------
Cheers,
Stan H.
----------